OpenStudy (anonymous):

square root of 3x-2 + square root of X+3 = 7

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\sqrt{3X-2}+\sqrt{X+3 = 7}\]

OpenStudy (anonymous):

=7 is not supposed to be under the radical

myininaya (myininaya):

get one square root by itself and then square both sides \[\sqrt{3x-2}=7-\sqrt{x+3}\] \[3x-2=7^2-2*7*\sqrt{x+3}+(x+3)\]

myininaya (myininaya):

\[3x-2=49-14\sqrt{x+3}+x+3\]

OpenStudy (anonymous):

4x+1=49 4x=48 x=12

myininaya (myininaya):

\[3x-2-x-3-19=-\sqrt{x+3}\]

myininaya (myininaya):

\[2x-24=-\sqrt{x+3}\]

OpenStudy (anonymous):

uhhh where did you get the 7*2 from in the first message?

myininaya (myininaya):

(a+b)^2=a^2+2*ab+b^2 (a-b)^2=a^2-2*ab+b^2

myininaya (myininaya):

now we square both sides again and time we get rid of all the square roots \[(2x-24)^2=(-\sqrt{x+3})^2\]

OpenStudy (anonymous):

ummm...i have no clue what that means...idk where u r pulling the 2 and 7 from to get 14

myininaya (myininaya):

\[(2x)^2-2*2x*24+24^2=x+3\]

myininaya (myininaya):

are you sure you don't know how to multiply (a-b)(a-b)?

OpenStudy (anonymous):

ummm maybe...idk what a-b is though...

myininaya (myininaya):

a minus b?

myininaya (myininaya):

a number - a different number? ?

OpenStudy (anonymous):

ok....but im not sure why that comes up in this equation or what you are subtracting...

myininaya (myininaya):

we have a-b on the left hand side i squared it to get rid of the square root but since i squared that side i squared the other side

OpenStudy (anonymous):

i understand that but how did you end up with a 14 before squaring it?

myininaya (myininaya):

14 after squareing

OpenStudy (anonymous):

ok....but what is 14? there's only a 7

myininaya (myininaya):

(a-b)^2=a^2-2*ab+b^2 a=7 b=sqrt(x+3) (7-sqrt(x+3))=7^2-2*7(sqrt(x+3))+(sqrt(x+3))^2

myininaya (myininaya):

that one side i lefted out a square (7-sqrt(x+3))^2*

OpenStudy (anonymous):

there is no -2+7 in the problem though

myininaya (myininaya):

no there isnt

OpenStudy (anonymous):

so then how r u squaring it?

myininaya (myininaya):

what do you mean? i'm scaring it (a-b)^2=a^2-2ab+b^2 do you now understand this part?

myininaya (myininaya):

squaring* not scaring lol

myininaya (myininaya):

(a-b)^2=(a-b)(a-b)=a^2-ab-ab+b^2=a^2-2ab+b^2 do you not agree with this?

OpenStudy (anonymous):

if i knew what that meant maybe i would...

myininaya (myininaya):

you kind of need to know how multiply binomials before you can do this type of problem

myininaya (myininaya):

would you like to practice multiplying binomials?

OpenStudy (anonymous):

no! i know how to multiple binomials... i just dont know what you are saying

myininaya (myininaya):

(a-b)^2=a^2-2ab+b^2 all i did was multiply a-b times a-b a-b is a binomial this is multiplying binomials you don't understand this so it seems you need to practice multiplying binomials

OpenStudy (anonymous):

but there are numbers that are not even in the problem

myininaya (myininaya):

a is 7 b is sqrt(x+3)

myininaya (myininaya):

=7^2-2*7*sqrt(x+3)+(sqrt(x+3))^2*